OpenStudy (anonymous):

A train crosses a pole in 15 seconds, and it crosses a 100 meters platform in 25 seconds, then what is the Length of the train.

OpenStudy (anonymous):

The pole has no real dimension so v=(100-0)/(25-15) m/s=10 m/s x=v*15=150 meters Not sure on my initial assumption about the pole.